OpenStudy (anonymous):

PLEASE HELP One number is four times another number. The larger number is also 87 more than the smaller number. Find the numbers. Choices for the smaller number 29, 30, 33. (STEP BY STEP TO HELP BETTER UNDERSTAND)

OpenStudy (kkutie7):

Ln= larger number Sn= smaller number Ln=4*Sn Ln=87+Sn you can use x and y if you prefer.

OpenStudy (anonymous):

lets do x and y okay?

OpenStudy (kkutie7):

ok sounds good so here. x= larger number y= smaller number x=4*y x=87+y

OpenStudy (anonymous):

okayyyy so whats our first step??

OpenStudy (kkutie7):

solve one of the equations and plug into the other. \[\frac{x}{4}=y\] \[x=87+(\frac{x}{4})\] now solve that

OpenStudy (anonymous):

sooo 87 goes over 4?

OpenStudy (kkutie7):

\[x-\frac{x}{4}=87\]

OpenStudy (anonymous):

what? im confused that doesn't make sense i thought we was adding ?

OpenStudy (kkutie7):

minor change. you want to deal with the x's first here is the next text \[\frac{4x}{4}-\frac{x}{4}=87\]

OpenStudy (kkutie7):

now subtract

OpenStudy (anonymous):

 like this ? i dont think im doing this right?!